OpenStudy (anonymous):

a+1/a=1 then the value of cube of a how tosolve it step by step

OpenStudy (shubhamsrg):

you mean a + (1/a) =1 ??

Parth (parthkohli):

@satishsheoran Some reply would be helpful for us to know what the question is :)

Parth (parthkohli):

\(\Rightarrow \Large {a^2 \over a} + {1 \over a} = 1 \) \(\Rightarrow\Large {a^2 + 1 \over a} = 1 \) \(\Rightarrow a^2 + 1 = a \) \(\Rightarrow a^2 - a + 1 = 0\) Solve.

OpenStudy (anonymous):

thnx